# Mechanical hysterons with tunable interactions of general sign

### v1

@@ Abstract
Discussion.— Here we have demonstrated a design for mechanical hysterons with continuously tunable thresholds and interactions, whose states can be controlled manually or by coupling to an external mechanical drive. Recent work has demonstrated other mechanical designs that can count modulo two [31], count driving cycles up to $n$ [30], or encode the magnitude of a mechanical force into a binary string [32]. Although those systems use bistable mechanical elements, each one uses a different strategy to navigate the pathways between its states. In contrast, we have presented a general platform that can be reconfigured to perform all of the above functions, plus a latching behavior that was recently identified in hysteron simulations [23]. Crucial to this versatility was the realization of cooperative and frustrated interactions that can be reciprocal or non-reciprocal.
There should of course be other ways to design physical hysterons with general interactions; Shohat and van Hecke [33] recently studied two-dimensional networks of bilinear hysterons as a route to exotic pathways, including mixed interactions and multiperiodic cycles. Understanding what are the general constraints on $\gamma_{i,0}^{\pm}$ and $J_{ij}$ for any realizable metamaterial is an open question.
One principle of our approach was to use nonlinearities sparingly, so that the behaviors of the system could be rationalized as simply as possible. We used linear springs and nearly-frictionless bearings in the experiment, and our analysis using small rotor angles and an affine relationship between torque and the global drive was able to capture the essential behaviors of two interacting hysterons (Fig. 3).
Designing the response of larger collections of mechanical hysterons presents some intriguing challenges. Although we report an exact kinematic model and a mapping to switching thresholds and interactions for small rotor angles, in practice we worked at relatively large rotor angles and proceeded through a combination of physical reasoning and trial-and-error. This approach might not work for larger hysteron machines. As an alternative, our kinematic model can provide the basis for computational inverse design methods, which could generate sets of parameters for a particular desired behavior. Another possibility for targeting behaviors in large systems is through supervised learning [34], which was recently implemented in experiments on spring networks [35]. Our system offers a wealth of possible learning degrees of freedom, as there are numerous geometric parameters that can be tuned continuously to affect the switching thresholds, including $x_{i}$, $\theta_{i}^{\pm}$, $\ell_{ij}$, and $L_{i}$. Understanding whether this approach can be used achieve a specific transition graph or a desired response under cyclic drive could be a fertile direction for future work.
@@ Acknowledgements.
I thank Nathan Keim for suggesting how to apply a global drive, for providing the parameters used in Fig. 6A, and for many fruitful discussions. I thank Vidyesh Anisetti, Eadin Block, Christian Santangelo, and Jennifer Schwarz for contributing to early hysteron designs, and I thank Chloe Lindeman, Dor Shohat, and Martin van Hecke for comments on the manuscript. This work was initiated at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452.
@@ Appendix A APPENDIX
Torques for general coupling.— We now outline the case where the two springs connecting hysterons $i$ and $j$ are not necessarily paired with equal spring constants or rest lengths, and where they may not have symmetrical mounting positions. We begin by considering a single coupling spring of stiffness $k$ and rest length $x_{0}$ that mounts to rotors $1$ and $2$ at positions $\ell_{1}$ and $\ell_{2}$ from the fulcrums of the rotors, respectively, where the signs of $\ell_{1}$, $\ell_{2}$ indicate the side of the fulcrum where they attach. In the limit of small $\theta_{1}$, $\theta_{2}$ this coupling spring exerts a torque on hysteron 1 equal to:

### v2

@@ Abstract
Discussion.— Here we have demonstrated a design for mechanical hysterons with continuously tunable thresholds and interactions, whose states can be controlled manually or by coupling to an external mechanical drive. Recent work has demonstrated other mechanical designs that can count modulo two [33], count driving cycles up to $n$ [32], or encode an analog mechanical input into a binary string [34]. Although those systems use bistable mechanical elements, each one uses a different strategy to navigate the pathways between its states. In contrast, we have presented a general platform that can be reconfigured to perform all of the above functions, plus a latching behavior that was recently identified in hysteron simulations [24]. Crucial to this versatility was the realization of cooperative and frustrated interactions that can be reciprocal or non-reciprocal.
One principle of our approach was to use nonlinearities sparingly, so that the behaviors of the system could be rationalized as simply as possible. We used linear springs and nearly-frictionless bearings in the experiment, and our analysis using small rotor angles and an affine relationship between torque and the global drive was able to capture the essential behaviors of two interacting hysterons (Fig. 3). Other designs for interacting mechanical hysterons are also being proposed; Shohat and van Hecke [35] used theory and simulations to explore two-dimensional networks of bilinear hysterons as a route to exotic pathways, including mixed interactions and multiperiodic cycles. Understanding what are the general constraints on $\gamma_{i,0}^{\pm}$ and $J_{ij}$ for any realizable metamaterial is an open question.
Designing the response of larger collections of mechanical hysterons presents another set of challenges. Although we report an exact kinematic model and a mapping to switching thresholds and interactions for small rotor angles, in practice we worked at relatively large rotor angles and proceeded through a combination of physical reasoning and trial-and-error. This approach might not work for larger hysteron machines. As an alternative, our kinematic model can provide the basis for computational inverse design methods, which could generate sets of parameters for a particular desired behavior. Another possibility for targeting behaviors in large systems is through supervised learning [36], which was recently implemented in experiments on spring networks [37]. Our system offers a wealth of possible learning degrees of freedom, as there are numerous geometric parameters that can be tuned continuously to affect the switching thresholds, including $x_{i}$, $\theta_{i}^{\pm}$, $\ell_{ij}$, and $L_{i}$. Understanding whether this approach can be used achieve a specific transition graph or a desired response under cyclic drive could be a fertile direction for future work.
@@ Acknowledgements.
I thank Nathan Keim for suggesting how to apply a global drive, for providing the parameters used in Fig. 6A, and for many fruitful discussions. I thank Vidyesh Anisetti, Eadin Block, Christian Santangelo, and Jennifer Schwarz for contributing to early hysteron designs, and I thank Martin van Hecke, Nathan Keim, Chloe Lindeman, and Dor Shohat for comments on the manuscript. This work was initiated at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452.
@@ Appendix A Appendix
Rotor hysteron construction.— We construct each rotor hysterons from two $1/2$”-diameter optical posts attached in a T configuration, which we mount in a radial steel ball bearing affixed to an optical table. We remove the bearing grease to further lower their friction to rotation. We use steel extension springs to couple the rotors to one another and to a 1/2”-diameter steel shaft that is mounted in two linear bearings. This 72”-long rod translates freely along the horizontal direction and provides the global mechanical drive, $\gamma$. We define $\gamma$ as the horizontal displacement of this rod in centimeters, which we measure with a precision of $0.1$ cm via a ruler affixed under the rod.
Spring stiffnesses.— We measure the spring stiffnesses from force-versus-displacement data that we obtain by anchoring the springs to a known mass on an electronic balance, and recording the force on the balance as the top of the spring is extended and then retracted (Fig. 1B). These measurements show that the springs behave linearly over the range of extensions used in the experiments.
Torques for general coupling.— Here we derive expressions for the torques in the small-$\theta$ model, in the general case where the two springs connecting hysterons $i$ and $j$ are not necessarily paired with equal spring constants or rest lengths, and where they may not have symmetrical mounting positions. We begin by considering a single coupling spring of stiffness $k$ and rest length $x_{0}$ that mounts to rotors $1$ and $2$ at positions $\ell_{1}$ and $\ell_{2}$ from the fulcrums of the rotors, respectively, where the signs of $\ell_{1}$, $\ell_{2}$ indicate the side of the fulcrum where they attach. In the limit of small $\theta_{1}$, $\theta_{2}$ this coupling spring exerts a torque on hysteron 1 equal to: